OpenStudy (anonymous):

Solve using the addition method. What is the y-value of the solution to the system? 2x - 3y = -1 3x + 4y = 24

OpenStudy (anonymous):

Solve using the addition method. What is the y-value of the solution to the system? 2x - 3y = -1 3x + 4y = 24 4 3 2 none of the above

OpenStudy (anonymous):

you have to multiply one or both of those equations by a number in order to get eliminate

OpenStudy (anonymous):

mult 1st one by 3: 6x - 9y = -3 mult 2nd one by -2: -6x + -8y = -48 now add them together -17y = -51 divide by -17 on both sides to determine value of y y = 3

OpenStudy (anonymous):

alright so how would you do this If you were to solve the system using the addition method, what would you multiply the first equation by in order to make the y's add out of this system? -2x + 3y = 6 7x - 6y = 8

OpenStudy (anonymous):

look at the Y's...what can be multiplied by one or both of them to cancel each other out

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

which one is to be multipled by -3?

OpenStudy (anonymous):

the y's

OpenStudy (anonymous):

which equation? You will multiply that number by EVERY term in THAT equation.

OpenStudy (anonymous):

the top

OpenStudy (anonymous):

how would it look

OpenStudy (anonymous):

oh, sorry - the question asks you for the first equation. Yes that is correct -3 will allow you to cancel out the term of Y

OpenStudy (anonymous):

right so is my final answer -2

OpenStudy (anonymous):

-2x + 3y = 6 7x - 6y = 8 mult 1st by 2 : -4x +6y = 12 orig 2nd : 7x-6y = 8 now add : 3x = 20 x = 20/3 you should multiply by 2 NOT -3

OpenStudy (anonymous):

right so is y 2??

OpenStudy (anonymous):

or 4

OpenStudy (anonymous):

The question asks - what # do you multiply 1st equation to cancel the Y -2x + 3y = 6 MULTIPLY BY 2 7x - 6y = 8 Remeber, in algebra - what you do to one term in an equation - you must do to ALL

OpenStudy (anonymous):

okay so what will y end up as

OpenStudy (anonymous):

?????????????

OpenStudy (anonymous):

Y is eliminated - it cancels out.... x = 20/3 now to solve for Y - go back to one of the original equation and plug in value of X - solve for Y

OpenStudy (anonymous):

The objective of this is to eliminate or cancel out a term in order to solve for the other one. This is a method in solving equations simultaneously

OpenStudy (anonymous):

so my answers were previously wrong 2 -2 4

OpenStudy (anonymous):

help